Determne the area under the standard normal arve that lies between (a) Ζ:-165 and Z-165 (b) Ζ--051 and Ζ:0 and (c)2 (a) The area that lies between Z-165 and Z . 165i (Round to four decimal places as needed) -171 and Z:-001
Answer #1

As a general rule, we want to express all the negative Z values as positive and right tail probabilities (that is Z>)  as left tail probabilities (that is Z< ). We will use the standard normal tables.

a) The area that lies between Z=-1.65 and Z=1.65 is calculated as

P(-1.65 < Z < 1.65) P(Z < 1.65)-P(Z <-1.65) -P(Z < 1.65)-P(Z > 1.65) due to the symmetry of normal distribution. P(Z <-1.65)-P(Z > 1.65) = P(Z < 1.65)-(1-P(Z < 1.65)) = 2 × P(Z < 1.65)-1 2 × (0.5 + 0.4505)-1 = 0.9010 using standard normal tables, get the area for z=1.65

ans: The area that lies between Z=-1.65 and Z=1.65 is 0.9010

b) The area that lies between Z=-0.51 and Z=0 is

P(-0.51<Z< 0)-P(Z< 0)- P(Z<-0.51) P(Z < 0)-P(Z > 0.51) due to the symmetry of normal distribution. P(Z <-0.51)-P(Z > 0.51) = P(Z < 0)-(1-P(Z < 0.51)) = 0.5-(1-(0.5 + 0.1950)) using standard normal tables, get the area for z=0.51 0.1950

ans: The area that lies between Z=-0.51 and Z=0 is 0.1950

c) The area that lies between Z=-1.71 and Z=-0.01 is

P(-1.71 < Z <-0.01) = P(Z <-0.01)-P(Z <-1.71) P(Z>0.01) P(Z1.71) due to the symmetry of normal distribution = (1-P(Z < 0,01)) _ (1-P(Z < 1.71)) = P( Z < 1.71)-P(Z < 0.01) = (0.5 + 0.4564)-(0.5 0.0040) using standard normal tables, get the area for z= 1.71 and z=0.01 0.4524

ans: The area that lies between Z=-1.71 and Z=-0.01 is 0.4524